Circular No. 7075

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 Y2 (LI) Weidong Li, Department of Astronomy, University of California at

Berkeley, reports his discovery of a comet in the course of the Lick Observatory Supernova Search (cf. IAUC 6627; with the participation of M. Papenkova, E. Halderson, M. Modjaz, T. Shefler, J. Y. King, R. R. Treffers and A. V. Filippenko). The object was found automatically by the Katzman Automatic Imaging Telescope in the field of NGC 1041, but it was immediately recognized as a comet by Li, who then used the equipment to make a deliberate confirmatory observation, as well as follow-up observations on the following night. Available CCD astrometry:

1998 UT α_{2000} δ_{2000} m_{1} Observer
Dec. 26.23575 2^{h}40^{m}26.^{s}37 −5^{o}25′56′′.4 16.2 Li
26.26623 2 40 26.47 −5 25 27.3 16.1  _{”}
27.31124 2 40 32.08 −5 08 39.7 15.3  _{”}
27.31584 2 40 32.12 −5 08 34.1 15.3  _{”}
27.32030 2 40 32.08 −5 08 29.5 15.5  _{”}
28.00613 2 40 37.04 −4 57 25.8 15.0 Kowalski
28.02951 2 40 37.28 −4 57 02.8  _{”}
28.04815 2 40 37.20 −4 56 43.1  _{”}
28.13387 2 40 37.85 −4 55 23.0 15.5 Bell
28.14006 2 40 37.87 −4 55 17.3  _{”}
28.16014 2 40 37.99 −4 54 57.8  _{”}

W.-d. Li (Lick Observatory). 0.8-m KAIT. Object diffuse. R. Kowalski (Zephyrhills). 0.18-m reflector. Coma perhaps 18′′ across. G. Bell and G. Hug (Eksridge). 0.3-m Schmidt-Cassegrain.

Parabolic orbital elements follow. The comet is likely of short period.

T = 1999 Feb. 1.854 TT  ω = 335.^{o}683
Ω = 87.914 }2000.0
q = 2.67161 AU  i = 26.412
1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 28 2^{h}40.^{m}61 − 4^{o}57.^{′}5 2.091 2.698 118.^{o}1 18.^{o}7 14.9
Jan. 2 2 41.51 − 3 35.7 2.134 2.691 114.2 19.5 14.9
7 2 42.98 − 2 12.5 2.180 2.685 110.4 20.1 15.0
12 2 45.01 − 0 48.4 2.229 2.681 106.5 20.6 15.0
17 2 47.60 + 0 36.1 2.281 2.677 102.8 21.0 15.1
22 2 50.72 + 2 00.5 2.336 2.674  99.1 21.3 15.1
1998 December 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden